Victor Haddad sold 8 pumpkins for $2.75 each,9 for $2.00 each, 24 for $1.50 each, and 15 for $1.00 each. He receives $0.35 for each pumpkin sold plus a $10.00 bonus if his sales total $75.00 or more. How much did he receive?
Answer by JulietG(1812)

8 @ 2.75 = 22
9 @ 2 = 18
24 @1.5 = 36
Sales total = 22 + 18 + 36 = 76, so he earned his bonus.
.35 x total number of pumpkins sold (8+9+24=41).
.35 * 41 = $14.35
Add the $10 bonus
$14.35 + $10 = $24.35